A company produces fruity drinks that contain a percentage of real fruit juice. Drink A contains 20% real fruit juice and Drink B contains 5% real fruit juice. The company used 56.5 liters of real fruit juice to make 80 more liters of Drink B than liters of Drink A. Write a system of equations that could be used to determine the number of liters of Drink A made and the number of liters of Drink B made. Define the variables that you use to write the system.

Answer by mananth(16946)


Drink A contains 20% real fruit juice. x liters
and Drink B contains 5% real fruit juice. y liters
y=x+80
20%x +5%y =56.5
20x+5y==5650
-x + y = 80.00


20x + 5 y = 5650.00 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 1
5.00 x -5.00 y = -400.00
20.00 x 5.00 y = 5650.00
Add the two equations
25.00 x = 5250.00
/ 25.00
x = 210.00
plug value of x in (1)
-1.00 x + 1.00 y = 80.00
-210.00 + 1.00 y = 80.00
1.00 y = 290.00
y = 290.00
Ans x = 210 liters Drink A
y = 290 liters Drink B